40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9294     Entry time: Fri Oct 25 21:28:49 2013
Author: Masayuki 
Type: Update 
Category: LSC 
Subject: REFL PDs spectrum 

 I measured the spectrum of the REFL165 output using AG4395A. As this entry we put the directional coupler between REFL165 output and demod board input, so I measure the signal from the coupler during the PRMI was locked.

 After measure REFL165, I also measured REFL55 output in order to make sure that the signal is not smaller than noise because of coupler. I terminated the couple output of coupler on the REFL165, and take signal from REFL55 output port directly. Both plots seems same except for around the resonant frequency of each PDs. From this plot we cannot say that the coupler reduce signal to spectrum analyser too much.

 After this measurement I reconnected the REFL165 to analyser and reconnected the REFL55 output to demod board.

Attachment 1: REFL.png  20 kB  Uploaded Fri Oct 25 22:31:38 2013  | Hide | Hide all
REFL.png
Attachment 2: REFLspe.zip  14 kB  Uploaded Sat Oct 26 21:20:53 2013
ELOG V3.1.3-